OpenStudy (anonymous):

Solve the system by elimination. x+5y-4z=-10 2x-y+5z=-9 2x-10y-5z=0

OpenStudy (anonymous):

x+5y-4z=-10 2x-y+5z=-9 ||eliminating z 13x + 21 y = -86 2x-y+5z=-9 2x-10y-5z=0 ||eliminating z 4x - 11y = -9 13x + 21 y = -86 21x = -105, x = -5 and y = -1 and z = 0